OpenStudy (anonymous):

How many real number solutions does the equation 7x^2-5x+1=0 have? please explain steps to solve

OpenStudy (anonymous):

No real solution because Discriminant i.e. d= b^2 - 4ac <0 in this case Because here b= -5 a= 7 and c=1 thus putting values, d= (-5)^2 - 4 (7)(1) = 25 - 28 = -3 < 0

OpenStudy (anonymous):

oh ok, can u help me with another prob?

OpenStudy (anonymous):

O.k. is this clear??? then post it....

OpenStudy (anonymous):

yeah I understand it now, A volleyball player spikes the ball from the height of 9 feet with a downward velocity of 80 feet per second. To the nearest hundredth of a second, how long does the other team have to prevent the ball from hitting the ground?

OpenStudy (anonymous):

this one confuses me

OpenStudy (anonymous):

It means you simply need to calculate the time after which it will strike the ground... use 2nd law of motion...

OpenStudy (anonymous):

is the formula that someone before gave me correct?

OpenStudy (anonymous):

I don't know what formula anyone gave you.... If you post here , I can check it now...

OpenStudy (anonymous):

Ya, its correct only...